In 1998, the womens world record for the 100-meter free-style in swimming was 54.73 seconds. The records time has been decreasing by approximately 0.33 seconds a year. The men's record was 49.36 seconds and has been decreasing at about 0.18 seconds a year. If both records continue to decrease at these same rates, in how many years will the men and women have the same record time?
Answer by mananth(16946)

In 1998, the womens world record for the 100-meter free-style in swimming was 54.73 seconds. The records time has been decreasing by approximately 0.33 seconds a year. The men's record was 49.36 seconds and has been decreasing at about 0.18 seconds a year. If both records continue to decrease at these same rates, in how many years will the men and women have the same record time?
...
1998 women 54.73
0.33 seconds/year reduction.
..
Men = 49.36
0.18 seconds /year reduction.
...
In 1998 the difference in record time = 54.73-49.36= 5.37 seconds.
...
catch up time = 0.33-0.18=0.15 seconds
..
5.37/0.15= 35.8 years
m.ananth@hotmail.ca